the length of a rectangle is 8 feet more than the width the perimeter of the rectangle is 76 feet, find the dimensions of the rectangle
Answer by lwsshak3(11628)

the length of a rectangle is 8 feet more than the width the perimeter of the rectangle is 76 feet, find the dimensions of the rectangle
***
let x=width
x+8=length
2(length+width)=perimeter
2(x+x+8)=76
2x+8=38
2x=30
x=15
x+8=23
dimensions of rectangle: 15 by 23 ft